Is the equation 4x-1=16 linear?
Found 2 solutions by elima, khwang:
Answer by elima(1433)

Is the equation 4x-1=16 linear?
No, a linear equation has 2 variables.
ex.:
x+1=y
=)

Answer by khwang(438)

Yes, any equation of 1st order is linear.
Geometricaly,it may be a line , a plane or a hyperplane.
e.g 2 x + 5 = 0
3x - 5y = 6
4x + 7y -z = 9
Forget another wrong answer.
Kenny